# Quantum enhancements and entropic constraints to Boundary Time Crystals as sensors of AC fields

### v1

@@ Abstract
We investigate the use of a boundary time crystals (BTCs) as sensors of AC fields. Boundary time crystals are non-equilibrium phases of matter in contact to an environment, for which a macroscopic fraction of the many-body system breaks the time translation symmetry. We find an enhanced sensitivity of the BTC when its spins are resonant with the applied AC field, as quantified by the quantum Fisher information (QFI). The QFI dynamics in this regime is shown to be captured by a relatively simple ansatz consisting of an initial power-law growth and late-time exponential decay. We study the scaling of the ansatz parameters with resources (encoding time and number of spins) and identify a moderate quantum enhancement in the sensor performance through comparison with classical QFI bounds. Investigating the precise source of this performance, we find that despite of its long coherence time and multipartite correlations (advantageous properties for quantum metrology), the entropic cost of the BTC (which grows indefinitely in the thermodynamic limit) hinders an optimal decoding of the AC field information. This result has implications for future candidates of quantum sensors in open system and we hope it will encourage future study into the role of entropy in quantum metrology.
@@ 1 Introduction
Time crystals display robust coherence and correlations making them appealing candidates for application in various quantum technologies. In particular there have been several recent publications investigating their use in quantum metrology [16, 17, 18, 19]. Generally, metrology seeks to optimally estimate a parameter through observing the statistics of a system (the sensor) affected by this parameter [20, 21, 22, 23]. This parameter could be, for example, some intrinsic system property or the strength of an external field incident on the sensor. For sensing AC external fields in particular time crystals have been seen to be effective candidates. In Ref. [16] they showed that the persistent oscillations of an FTC allowed significant quantum enhancement in the sensing of an AC magnetic field incident on the system.
@@ 2 Model
where $\{\lambda_{i}\}$ are the eigenvalues of $\hat{\rho}_{g}$. Here we have suppressed the time dependence for conciseness.
Computing $\partial_{g}\hat{\rho}_{g}$ can be done numerically by either a finite difference in terms of $g$ or, as we shall show, by computing a second master equation in conjunction with Eq. (4). We can find the time evolution of $\partial_{g}\hat{\rho}_{g}$ by simply taking the time derivative of Eq. (4), this given by
$\partial_{t}\partial_{g}\hat{\rho}_{g}(t)=\hat{\mathcal{L}}_{B}[\partial_{g}\hat{\rho}_{g}(t)]-\\ -i[\hat{H}_{\text{ac}}(t),\hat{\rho}_{g}(t)+g\partial_{g}\hat{\rho}_{g}(t)].$ (9)
@@ 3 QFI Dynamics
In this section, we study the scaling of the QFI as a function of $N$ and $t$. We use its scaling behaviour to characterize the performance of the sensor and compare this performance with the limits of a classical sensor as well as the fundamental “Heisenberg” quantum limit. These limits are expressed in terms of the total resources used in the sensing protocol. Incorrect characterization of the resources used or omitting a certain resource, such as sensing time, can lead to results which seemingly surpass even the Heisenberg limit. We shall be careful to state which resources we consider in each result. All plots results generated here correspond to an initial state for which $\langle\hat{S}^{z}\rangle=S$.
@@ 5 Conclusions
Our hope is that this work encourages a deeper study into the role of entropy in quantum metrology.
The authors gratefully acknowledge the computing time granted on the supercomputer MOGON 2 at Johannes Gutenberg-University Mainz (hpc.uni-mainz.de). F.I. acknowledges financial support from Alexander von Humboldt foundation and the Brazilian funding agencies CAPES, CNPQ, and FAPERJ (Grants No. 308205/2019-7, No. E-26/211.318/2019, No. 151064/2022-9, and No. E-26/201.365/2022) and by the Serrapilheira Institute (Grant No. Serra 2211-42166). A.S. acknowledges financial support from the Spanish Agencia Estatal de Investigacion, Grant No. PID2022-141283NB-I00, the European Commission QuantERA grant ExTRaQT (Spanish MICINN project PCI2022-132965), by Ministerio de Ciencia e Innovación with funding from European Union NextGenerationEU(PRTR-C17.I1) and by Generalitat de Catalunya and the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call – Quantum Spain project, and the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2025 Agend. R.F. ackowledges financial support from PNRR MUR project PE0000023-NQSTI and by the European Union (ERC, RAVE, 101053159). D.G. and J.M. acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”), Project-ID 429529648 (project D04) and in part by the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 (“QuSiED”). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

# Boundary Time Crystals as AC sensors: enhancements and constraints

@@ Abstract
We investigate the use of a boundary time crystals (BTCs) as quantum sensors of AC fields. Boundary time crystals are non-equilibrium phases of matter in contact to an environment, for which a macroscopic fraction of the many-body system breaks the time translation symmetry. We find an enhanced sensitivity of the BTC when its spins are resonant with the applied AC field, as quantified by the quantum Fisher information (QFI). The QFI dynamics in this regime is shown to be captured by a relatively simple ansatz consisting of an initial power-law growth and late-time exponential decay. We study the scaling of the ansatz parameters with resources (encoding time and number of spins) and identify a moderate quantum enhancement in the sensor performance through comparison with classical QFI bounds. Investigating the precise source of this performance, we find that despite of its long coherence time and multipartite correlations (advantageous properties for quantum metrology), the entropic cost of the BTC (which grows indefinitely in the thermodynamic limit) hinders an optimal decoding of the AC field information. This result has implications for future candidates of quantum sensors in open system and we hope it will encourage future study into the role of entropy in quantum metrology.
@@ 1 Introduction
Time crystals display robust coherence and correlations making them appealing candidates for application in various quantum technologies. In particular there have been several recent publications investigating their use in quantum metrology [16, 17, 18, 19, 20]. Generally, metrology seeks to optimally estimate a parameter through observing the statistics of a system (the sensor) affected by this parameter [21, 22, 23, 24]. This parameter could be, for example, some intrinsic system property or the strength of an external field incident on the sensor. For sensing AC external fields in particular time crystals have been seen to be effective candidates. In Ref. [16] they showed that the persistent oscillations of an FTC allowed significant quantum enhancement in the sensing of an AC magnetic field incident on the system.
@@ 2 Model
where $\{\lambda_{i}\}$ are the eigenvalues of $\hat{\rho}_{g}$. Here we have suppressed the time dependence for conciseness. Computing $\partial_{g}\hat{\rho}_{g}$ can be done numerically by either a finite difference in terms of $g$ or, as we shall show, by computing a second master equation in conjunction with Eq. (2). We can find the time evolution of $\partial_{g}\hat{\rho}_{g}$ by simply taking the time derivative of Eq. (2), this given by
$\partial_{t}\partial_{g}\hat{\rho}_{g}(t)=\hat{\mathcal{L}}_{B}[\partial_{g}\hat{\rho}_{g}(t)]-i[\hat{H}_{\text{ac}}(t),\hat{\rho}_{g}(t)+g\partial_{g}\hat{\rho}_{g}(t)].$ (7)
@@ 3 QFI Dynamics
In this section, we study the scaling of the QFI as a function of $N$ and $t$. We use its scaling behaviour to characterize the performance of the sensor and compare this performance with the limits of a classical sensor as well as the fundamental “Heisenberg” quantum limit. These limits are expressed in terms of the total resources used in the sensing protocol. Incorrect characterization of the resources used or omitting a certain resource, such as sensing time, can lead to results which seemingly surpass even the Heisenberg limit. We shall be careful to state which resources we consider in each result. All results presented here correspond to an initial state for which $\langle\hat{S}^{z}\rangle=S$, they were were generated using QuantumOptics.jl [38] and code adapted from this framework.
@@ 5 Conclusions
In future it would be interesting to investigate the performance of the BTC as an AC sensor in the vicinity of its critical point; there are many examples of enhanced sensitivity near criticality [19, 45, 46, 47, 48, 49] and evidence that the associated scaling is robust to periodic driving [50]. Beyond this, we hope that our work encourages a deeper study into the role of entropy in quantum metrology.
Acknowledgements
The authors gratefully acknowledge the computing time granted on the supercomputer MOGON 2 at Johannes Gutenberg-University Mainz (hpc.uni-mainz.de).
Funding information
F.I. acknowledges financial support from Alexander von Humboldt foundation and the Brazilian funding agencies CAPES, CNPQ, and FAPERJ (Grants No. 308205/2019-7, No. E-26/211.318/2019, No. 151064/2022-9, and No. E-26/201.365/2022) and by the Serrapilheira Institute (Grant No. Serra 2211-42166). A.S. acknowledges financial support from the Spanish Agencia Estatal de Investigacion, Grant No. PID2022-141283NB-I00, the European Commission QuantERA grant ExTRaQT (Spanish MICINN project PCI2022-132965), by Ministerio de Ciencia e Innovación with funding from European Union NextGenerationEU(PRTR-C17.I1) and by Generalitat de Catalunya and the Spanish Ministry of Economic Affairs and Digital Transformation through the QUANTUM ENIA project call – Quantum Spain project, and the European Union through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2025 Agend. R.F. ackowledges financial support from PNRR MUR project PE0000023-NQSTI and by the European Union (ERC, RAVE, 101053159). D.G. and J.M. acknowledge financial support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via TRR 306 QuCoLiMa (“Quantum Cooperativity of Light and Matter”), Project-ID 429529648 (project D04) and in part by the QuantERA II Programme that has received funding from the European Union’s Horizon 2020 research and innovation programme under Grant Agreement No 101017733 (“QuSiED”). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.